for two equally likely, exhaustive and independent events a and b. what is the probability of a and b
1
Expert's answer
2019-04-16T10:02:05-0400

Suppose a and b are two events associated with a random experiment. The probability of union of two events is given by:


"P(a\cup b)=P(a)+P(b)-P(a\cap b)"

Equally likely events are events that have the same theoretical probability of occurring.


"P(a)=P(b)"

If a and b are independent events, the probability of both the events happening together


"P(a\cap b)=P(a)P(b)"

If a and b are exhaustive events, the probability of union of two events is


"P(a\cup b)=1"

Hence


"1=P(a)+P(a)-(P(a))^2"

"P(a)(2-P(a))=1"

If "P(a)=0", then "P(b)=P(a)=0" and "P(a\cup b)=0". False.

Therefore, "P(a) =\not 0" and


"2-P(a)=1"

"P(a)=1=P(b)"

"P(a\cap b)=(1)^2=1"